A LASER INDUCED FLUORESCENCE STUDY OF JET-COOLED HCCl IN THE REGION 515-615 NM

\underline{H. FAN}, I. IONESCU, C. ANNESLEY, AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

We report new measurements of the laser induced fluorescence spectra of the HCCl A^{1}A′′-X^{1}A′system in the 515-615 nm region. The spectra were obtained under jet-cooled conditions using a pulsed discharge source. The pure bending transitions 2_{0}^{n} with n=5-8 and combination bands 2_{0}^{n}3_{0}^{m} with n=4-6 and m=1,2 have been measured and the rotational structure fit to a asymmetric top Hamiltonian, which yielded precise values for the band origins and effective rotational constants. Consistent with previous work, only sub-bands with K′_{a} = 0 appear strongly in the spectra. Fluorescence lifetimes were measured for rotational levels in each band, and the energy dependence of the lifetimes will be discussed.